### v1

# Vertex ordering characterizations of interval $r$- graphs

Abstract
An $r$-partite graph is an interval $r$-graph if corresponding to each vertex we can assign an interval of the real line such that two vertices $u$ and $v$ of different partite sets are adjacent if and only if their corresponding intervals intersect. In this paper, we provide two vertex-ordering characterizations of interval $r$-graphs and identify forbidden patterns for interval $r$-graphs in terms of specific orderings of their vertices.
Keywords: interval $r$-graphs, vertex ordering, generalized interval ordering, $r$-interval ordering , forbidden patterns
1 Introduction
A graph $G=(V,E)$ is called an interval graph, if there exists a family $\mathcal{I}=\{I_{v}:v\in V\}$, of intervals of the real line such that, for $a,b\in V$ the vertices $a$ and $b$ are adjacent in $G$ if and only if $I_{a}\cap I_{b}\neq\emptyset$. A bigraph $B=(X,Y,E)$ is said to be an interval bigraph if there exists a family $\mathcal{I}=\{I_{v}:v\in X\cup Y\}$ of intervals of the real line such that for all $x\in X$ and $y\in Y$ the vertices $x$ and $y$ are adjacent in $B$ if and only if $I_{x}\cap I_{y}\neq\emptyset$. Interval bigraphs were introduced in [5] and have been studied in [2, 3, 6, 7]. Interval digraphs were introduced by Sen et al. [9]. It was shown in [3] that the notions of interval digraphs and interval bigraphs are equivalent. Interval bigraphs or interval digraphs have become of interest in such new areas as graph homomorphism, e.g [4]. The concept of interval bigraph was generalized by Brown [1]. A graph $G=(V,E)$ is called interval $r$-graph if the vertex set $V$ can be partitioned into $r$ stable sets say $V_{1}$, $V_{2}$,…, $V_{r}$ and corresponding to each vertex we can assign an interval such that two vertices $u$ and $v$ are adjacent in $G$ if and only if the corresponding intervals $I_{u}$ and $I_{v}$ have non empty intersection and $u$ and $v$ belong to different partite set of vertices.
In this paper, we study different types of vertex orderings of interval $r$-graphs and, with the help of these orderings, we characterize the class of interval $r$-graphs. Finally, we provide forbidden patterns for interval $r$-graphs with respect to a particular vertex ordering.
2 Main Result
In this section, we introduce several types of vertex orderings for $r$-partite graphs and using these orderings we characterize interval $r$-graphs. Based on these orderings, we provide characterizations and identify forbidden patterns for interval $r$-graphs.
We begin with the definition of generalized interval ordering of the vertices of an $r$-partite graph.
Definition 1.
Let $G=(X_{1},X_{2},\ldots,X_{r},E)$ be an $r$-partite graph of order $n$. Arrange the vertices of $G$ linearly and label them $v_{1},v_{2},\ldots,v_{n}$ such that the vertex $v_{i}$ is placed on the $i$-th positive integer of the real line. Let $X=\bigcup_{i=1}^{r}X_{i}$ denote the vertex set of $G$. The set $X$ is said to admit a generalized interval ordering if the following condition is satisfied:
  • For any edge $v_{i}v_{j}\in E$ with $i>j$, where $v_{i}\in X_{\alpha}$, $v_{j}\in X_{\beta}$, and $\alpha\neq\beta$, it holds that $v_{\ell}v_{j}\in E\quad\text{for all }v_{\ell}\notin X_{\beta}\text{ with }\ell\in\{j+1,j+2,\ldots,i-1\}.$
We shall now characterize interval $r$-graphs using this ordering.
Theorem 1.
An $r$-partite graph $G=(X_{1},X_{2},...X_{r},E)$ is a interval $r$-graph if and only if the vertex set $X=\bigcup_{i=1}^{r}X_{i}$ of $G$ admits a generalized interval ordering.
Proof.
Necessity: Let $G=(X_{1},X_{2},\dots,X_{r},E)$ be an interval $r$-graph. Then there exists a class $\mathcal{I}=\{I_{v}:v\in\bigcup_{i=1}^{r}X_{i}\}$ of intervals corresponding to each vertex $v$ of $G$, such that $uv\in E$ if and only if $I_{u}\cap I_{v}\neq\emptyset$, where $u$ and $v$ belong to different partite sets. Without loss of generality, we assume that all intervals have distinct endpoints.
Now, order the vertices of $G$ as $v_{1},v_{2},\dots,v_{n}$ according to the increasing order of the left endpoints of their corresponding intervals, where $n$ denotes the order of the $r$-partite graph $G$. We will show that this ordering is a generalized interval ordering for the $r$-partite graph $G$.
Suppose $v_{i}v_{j}\in E$, where $i>j$ (and $v_{i}\in X_{\alpha}$, $v_{j}\in X_{\beta}$, $\alpha\neq\beta$). According to our ordering, the left endpoint of $v_{i}$ must lie within the interval $I_{v_{j}}$ corresponding to $v_{j}$. From Figure 1, it follows that the left endpoints of the intervals $I_{v_{j+1}},I_{v_{j+2}},I_{v_{j+3}},\dots,I_{v_{i-1}}$ are all contained in $I_{v_{j}}$. Thus $v_{l}v_{j}\in E$ if $v_{l}\notin X_{\beta}$ for $l\in\{j+1,\dots,i-1\}$. Hence, by Definition 1, we conclude that the ordering $v_{1},v_{2},\dots,v_{n}$ of the vertices of $G$ is a generalized interval ordering.
$v_{1}$$v_{2}$$v_{j}$$v_{j+1}$$v_{j+2}$$v_{i-1}$$v_{i}$$v_{i+1}$$v_{n-1}$$v_{n}$$I_{v_{i}}$$I_{v_{j}}$
Figure 1: Left end point of $I_{v_{i}}$ is contained in $I_{v_{j}}$.
Sufficiency: Let $G=(X_{1},X_{2},\dots,X_{r},E)$ be an $r$-partite graph and the ordering $v_{1},v_{2},\dots,v_{n}$ of its vertices is a generalized interval ordering.
Now, let $m_{i}=\max\left(\{\,j:v_{i}v_{j}\in E\,\}\cup\{i\}\right)$. Construct an interval $I^{\prime}_{v_{i}}$ for each vertex $v_{i}$ $(1\leq i\leq n)$ as $I^{\prime}_{v_{i}}=\left[\,i,m_{i}+\frac{1}{2}\,\right]$. First we shall show that the collection $\mathcal{I}^{\prime}=\{\,I^{\prime}_{v_{i}}:1\leq i\leq n\,\}$ of intervals satisfies the necessary conditions for $G$ to be an interval $r$-graph.
Let $v_{i}$ and $v_{j}$ be two vertices belonging to different partite sets of $G$ such that $v_{i}v_{j}\in E$ and $i>j$. According to our construction, $I^{\prime}_{v_{i}}=\left[\,i,m_{i}+\frac{1}{2}\,\right]\quad\text{and}\quad I^{\prime}_{v_{j}}=\left[\,j,m_{j}+\frac{1}{2}\,\right]$. Since $v_{i}v_{j}\in E$ and $i>j$, we have $m_{j}\geq i$, which implies $m_{j}+\frac{1}{2}\geq i$. Therefore, the interval $I^{\prime}_{v_{j}}$ contains $i$, and hence $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{j}}\neq\emptyset$.
Conversely, let $v_{i}$ and $v_{j}$ be two vertices of $G$ from different partite sets such that $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{j}}\neq\emptyset$ and $i>j$. Then $i\in I^{\prime}_{v_{j}}$, which implies $m_{j}\geq i$. Thus, $j<i\leq m_{j}$, and since $v_{j}v_{m_{j}}\in E$, by the definition of a generalized interval ordering we have $v_{i}v_{j}\in E$. Hence, $v_{i}v_{j}\in E\quad\Longleftrightarrow\quad I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{j}}\neq\emptyset$. Therefore, $G$ is an interval $r$-graph.
Before presenting another vertex-ordering characterization, we define the concept of almost consecutive ones in the rows and columns of the adjacency matrix of $r$-partite graphs.
Definition 2.
Let $G=(X_{1},X_{2},\ldots,X_{r},E)$ be an $r$-partite graph and let A be the adjacency matrix of $G$. A row (say, the $i$-th row) is said to have almost consecutive ones if, between any two ones in that row, whenever a zero appears at the position $(v_{i},v_{k})$, the vertices $v_{i}$ and $v_{k}$ belong to the same partite set of $G$.
Definition 3.
Let $G=(X_{1},X_{2},\ldots,X_{r},E)$ be an $r$-partite graph and let A be the adjacency matrix of $G$. A column (say, the $j$-th column) is said to have almost consecutive ones if, between any two ones in that column, whenever a zero appears at the position $(v_{l},v_{j})$, the vertices $v_{l}$ and $v_{j}$ belong to the same partite set of $G$.
We now introduce another type of vertex ordering for $r$-partite graphs, referred to as the $r$-interval ordering. Using this ordering, we characterize the class of interval $r$-graphs.
Let $G=(X_{1},X_{2},\dots,X_{r},E)$ be an $r$-partite graph of order $n$, and let $v_{1},v_{2},\dots,v_{n}$ be an ordering of its vertices. Place these vertices on the real line such that the $i$-th vertex $v_{i}$ is placed on the positive integer point $i$.
Let A denote the adjacency matrix of $G$, where both rows and columns are arranged in the increasing order of vertex indices. Consider any row of A (say, the $i$-th row). Define $R_{i}$ as the set of entries of $1$ in this row that appear almost consecutively, beginning with the column corresponding to the vertex $v_{s_{i}}$, where $v_{s_{i}}$ is the first vertex to the right of $v_{i}$ that belongs to a different partite set and $s_{i}>i$. If no $v_{i}v_{s_{i}}\in E$, then $R_{i}=\emptyset$. Otherwise, the sequence continues rightward until the last $1$ in this almost consecutive manner is reached.
Similarly, consider any column of A (say, the $j$-th column). Define $C_{j}$ as the set of entries of $1$ in this column that appear almost consecutively, beginning with the row corresponding to the vertex $v_{s_{j}}$, where $v_{s_{j}}$ is the first vertex to the right of $v_{j}$ that belongs to a different partite set and $s_{j}>j$. If no $v_{j}v_{s_{j}}\in E$, then $C_{j}=\emptyset$. Otherwise, the sequence continues downward until the last $1$ in this almost consecutive manner is reached.
An ordering $v_{1},v_{2},\dots,v_{n}$ of the vertices of $G$ is called an $r$-interval ordering if the sets $R_{i}$ and $C_{j}$ together contain all the $1$’s in the adjacency matrix A.
Theorem 2.
An $r$-partite graph $G=(X_{1},X_{2},\dots,X_{r},E)$ is an interval $r$-graph if and only if its vertex set $X=\bigcup_{i=1}^{r}X_{i}$ admits an $r$-interval ordering.
Proof.
Necessity: Let $G=(X_{1},X_{2},\ldots,X_{r},E)$ be an interval $r$-graph of order $n$. Then there exists an interval model
$\mathcal{I}=\{I_{v}:v\in X=\bigcup_{i=1}^{r}X_{i}\}$
such that for any two vertices $u,v$ belonging to different partite sets,
$uv\in E\quad\text{if and only if}\quad I_{u}\cap I_{v}\neq\emptyset.$
Without loss of generality, we may assume that the interval model $\mathcal{I}$ satisfies the following conditions:
  1. 1.
    All intervals are closed, i.e., each interval includes its endpoints.
  2. 2.
    No two intervals share the same left endpoint.
Label the vertices of $G$ as $v_{1},v_{2},\ldots,v_{n}$ according to the increasing order of the left endpoints of their corresponding intervals. Arrange both the rows and columns of the adjacency matrix $A$ of $G$ according to this vertex ordering.
We claim that, under this arrangement, the sets $R_{i}$ and the sets $C_{j}$ together contain all the 1’s of the adjacency matrix $A$.
Suppose an entry of $A$ is $a_{ij}=1$. This implies that $v_{i}$ and $v_{j}$ are adjacent in $G$, i.e., $v_{i}v_{j}\in E$. Let $v_{i}\in X_{\alpha}$ and $v_{j}\in X_{\beta}$, where $\alpha\neq\beta$. We distinguish two cases according to the relative positions of $v_{i}$ and $v_{j}$ in the vertex ordering.
Case 1: $i>j$.
From the interval representation (see Figure 2), any vertex appearing between $v_{j}$ and $v_{i}$ in this ordering, and belonging to a partite set other than $X_{\beta}$, will be adjacent to $v_{j}$. Consequently, $C_{j}$ will contain the 1 corresponding to the entry $(i,j)$ in the adjacency matrix.
$v_{1}$$v_{2}$$v_{j}$$v_{j+1}$$v_{j+2}$$v_{i-1}$$v_{i}$$v_{i+1}$$v_{n-1}$$v_{n}$$I_{v_{i}}$$I_{v_{j}}$
Figure 2: $I_{v_{i}}\cap I_{v_{j}}\neq\emptyset$, where $i>j$.
Case 2: $i<j$.
Similarly, any vertex appearing between $v_{i}$ and $v_{j}$ that does not belong to the same partite set $X_{\alpha}$ will be adjacent to $v_{i}$. Hence, $R_{i}$ will contain the 1 corresponding to the entry $(i,j)$ in the adjacency matrix.
$v_{1}$$v_{2}$$v_{i}$$v_{i+1}$$v_{i+2}$$v_{j-1}$$v_{j}$$v_{j+1}$$v_{n-1}$$v_{n}$$I_{v_{j}}$$I_{v_{i}}$
Figure 3: $I_{v_{i}}\cap I_{v_{j}}\neq\emptyset$, where $i<j$.
In either case, the 1 at the $(i,j)$-th entry is contained in one of $R_{i}$ or $C_{j}$. Therefore, the sets $R_{i}$ and $C_{j}$ together include all the 1’s of the adjacency matrix $A$.
Sufficiency: Consider an $r$-partite graph $G=(X_{1},X_{2},\ldots,X_{r},E)$, where the vertex set $X=\bigcup_{i=1}^{r}X_{i}$ is ordered as $v_{1},v_{2},\ldots,v_{n}$. Assume that this ordering is an $r$-interval ordering.
Assume if the rows and columns of the adjacency matrix of $G$ are arranged according to this ordering, then the sets $R_{i}$ and $C_{j}$ together contain all the 1’s in the adjacency matrix.
We now construct an interval model for $G$. If $R_{i}=\emptyset$, define $I_{v_{i}}=[i,i]$. Otherwise, let $r_{i}$ be the greatest column index such that the entry at the $(i,r_{i})$-th position of $A$ is 1. Then define $I_{v_{i}}=[i,r_{i}]$.
It remains to show that the family of intervals $\mathcal{I}=\{I_{v_{i}}:1\leq i\leq n\}$ serves as an interval model for the $r$-partite graph $G$.
Suppose $I_{v_{i}}\cap I_{v_{j}}\neq\emptyset$, where ($i<j$) and $v_{i}$, $v_{j}$ belong to different partite sets. Then by the construction of the intervals we can say that $j\in I_{v_{i}}$ $\implies\mathcal{R}_{i}$ contains the $1$ at position $(i,j)$ of the adjacency matrix of $G$. Therefore $v_{i}$ is adjacent to $v_{j}$.
Conversely, let $v_{i}$ is adjacent to $v_{j}$ in $G$, where $(i>j)$ , say. Therefore $\mathcal{R}_{j}$ will contain the 1 at the $(j,i)$-th entry. Implies $I_{v_{j}}$ contains $i$, hence by the definition of $I_{v_{i}}$ we have $I_{v_{i}}\cap I_{v_{j}}\neq\emptyset$. Therefore the class of intervals as constructed above is an interval model for the $r$-partite graph and so $G=(X_{1},X_{2},...,X_{r},E)$ is an interval $r$-graph.
Example 1. Consider the following $3$-partite graph. Using the above theorem we show that it is an interval $3$-graph.
$v_{7}$$v_{5}$$v_{3}$$v_{4}$$v_{6}$$v_{10}$$v_{1}$$v_{2}$$v_{9}$$v_{8}$
Figure 4: A 33-partite graph.
[…]
Figure 5: The adjacency matrix of the 33 partite graph in Figure 2, where the rows and columns are arranged according to the increasing order of their indices, and corresponding $\mathcal{R}_{i}$’s and $\mathcal{C}_{j}$’s are shown in the elliptic regions. And its interval representation is as follows:
Red: $I_{v_{1}}=[1,3]$, $I_{v_{6}}=[6,9]$, $I_{v_{10}}=[10,10]$
Green: $I_{v_{2}}=[3,4]$, $I_{v_{8}}=[8,10]$, $I_{v_{9}}=[9,9]$
Black: $I_{v_{3}}=[3,3]$, $I_{v_{4}}=[4,6]$, $I_{v_{5}}=[5,8]$, $I_{v_{7}}=[7,9]$.
Note that the zeros inside the elliptic regions of $\mathcal{R}_{5}$ and $\mathcal{C}_{5}$, as shown in Figure 3, do not belong to the sets $\mathcal{R}_{5}$ and $\mathcal{C}_{5}$. They simply indicate that the 1’s in $\mathcal{R}_{5}$ and $\mathcal{C}_{5}$ are not strictly consecutive, but rather almost consecutive.
Pavol Hell and Jing Huang [6] characterized interval bigraphs using forbidden patterns with respect to a specific vertex ordering in the following theorem.
Theorem 3 ([6]).
Let $H$ be a bipartite graph with bipartition $(X,Y)$. Then the following statements are equivalent:
  • $H$ is an interval bigraph;
  • the vertices of $H$ can be ordered $v_{1}$, $v_{2}$, …, $v_{n}$, so that there do not exist $a<b<c$ in the configurations in Figure 6. (Black vertices are in $X$, red vertices in $Y$, or conversely, and all edges not shown are absent.)
  • the vertices of $H$ can be ordered $v_{1}$, $v_{2}$, …, $v_{n}$, so that there do not exist $a<b<c<d$ in the configurations in Figure 7.
$v_{a}$$v_{b}$$v_{c}$
Figure 6: Forbidden pattern.
[…]
Figure 7: Forbidden patterns.
It is a natural question whether the classes of interval $r$-graphs ($r\geq 3$) can be characterized by a finite collection of forbidden patterns with respect to some specific ordering of their vertices. The following theorem provides an answer to this question.
Theorem 4.
Let $G=(X_{1},X_{2},\ldots,X_{r},E)$ be an $r$-partite graph ($r\geq 3$). Then the following statements are equivalent:
  1. i.
    $G$ is an interval $r$-graph;
  2. ii.
    The vertices of $G$ can be ordered $v_{1},v_{2},\ldots,v_{n}$ such that no indices $i<j<k$ occur in the configurations shown in Figure 8. (Here, different colors indicate that vertices belong to different partite sets; moreover, all edges not explicitly drawn are absent.)
[…]
Figure 8: Forbidden patterns.
Proof.
Necessity. Let $G=(X_{1},X_{2},\ldots,X_{r},E)$ $(r\geq 3)$ be an interval $r$-graph. Let $\mathcal{I}=\{I_{v}:v\in\bigcup_{i=1}^{r}X_{i}\}$ denote an interval model of $G$.
Without loss of generality, we may assume that the interval model $\mathcal{I}$ satisfies the following conditions:
  1. 1.
    All intervals are closed, i.e., each interval includes its endpoints.
  2. 2.
    No two intervals share the same left endpoint.
Arrange the vertices of $G$ according to the increasing order of the left endpoints of the corresponding intervals.
Consider any three vertices $v_{i},v_{j},v_{k}$ of $G$ belonging to three distinct partite sets $X_{\alpha},X_{\beta},X_{\gamma}$, respectively, where $i<j<k$. Suppose that $v_{i}$ and $v_{k}$ are adjacent. According to the vertex ordering, the left endpoint of the interval $I_{v_{k}}$ must lie within the interval $I_{v_{i}}$. Consequently, the left end points of the intervals $I_{v_{i+1}},I_{v_{i+2}},\ldots,I_{v_{k-1}}$ are contained in $I_{v_{i}}$, and therefore each of them intersects $I_{v_{i}}$.
It follows that every vertex indexed between $v_{i}$ and $v_{k}$ that does not belong to $X_{\alpha}$ must be adjacent to $v_{i}$. In particular, $v_{j}\in X_{\beta}$ is adjacent to $v_{i}$. Hence, the configurations (ii) and (iii) shown in Figure 8, cannot occur in an interval $r$-graph.
Similarly, if the vertices $v_{j}$ and $v_{k}$ belong to the same partite set, then the configuration (i) illustrated in Figure 8, also cannot occur in an interval $r$-graph. Thus, none of the three configurations shown in Figure 8 can occur in $G$.
Sufficiency: Suppose $G=(X_{1},X_{2},\ldots,X_{r},E)$ is an $r$-partite graph that admits an ordering $v_{1},v_{2},\ldots,v_{n}$ such that none of the configurations illustrated in Figure 8 occur. We shall show that $G$ is an interval $r$-graph.
Place the vertices of $G$ on the real line such that each vertex $v_{i}$ is positioned at the positive integer point $i$, for $1\leq i\leq n$. We now construct an interval corresponding to each vertex $v_{i}$ ($1\leq i\leq n$).
For a vertex $v_{i}\in X_{\alpha}$, define $t_{i}$ to be the largest vertex index (in the rightward direction from $v_{i}$) corresponding to a vertex $v_{t_{i}}\in X\setminus X_{\alpha}$ such that all vertices $v_{l}\in X\setminus X_{\alpha}$, where $i<l\leq t_{i}$, are adjacent to $v_{i}$. If no such vertex index exists, then set $t_{i}=i$.
Now define a family of intervals
$\mathcal{I}^{\prime}=\{I^{\prime}_{v_{i}}:1\leq i\leq n\}$
corresponding to the vertex set of $G$, where
$I^{\prime}_{v_{i}}=[i,\,t_{i}+\tfrac{1}{2}].$
We will show that this family of intervals is an interval representation of the graph $G$.
Claim 1. If two vertices $v_{i}\in X_{\alpha}$ and $v_{k}\in X_{\gamma}$ (with $\alpha\neq\gamma$ and $i<k$) satisfy $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{k}}\neq\emptyset$, then $v_{i}$ and $v_{k}$ are adjacent in $G$.
Proof. From the definition of the intervals,
$I^{\prime}_{v_{i}}=[i,t_{i}+\tfrac{1}{2}],\quad I^{\prime}_{v_{k}}=[k,t_{k}+\tfrac{1}{2}].$
If $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{k}}\neq\emptyset$, then $k\in I^{\prime}_{v_{i}}$, implying
$i\leq k\leq t_{i}+\tfrac{1}{2}\implies i\leq k\leq t_{i}.$
Hence, $v_{i}$ is adjacent to $v_{k}$.
Claim 2. Conversely, suppose $v_{i}$ and $v_{k}$ ($i<k$) are adjacent in $G$, where $v_{i}\in X_{\alpha}$ and $v_{k}\in X_{\gamma}$. We will show that $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{k}}\neq\emptyset$.
Proof. For contradiction, assume that $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{k}}=\emptyset$. Then, by definition of the intervals, $t_{i}<k$. This means that $v_{i}$ is not adjacent to every vertex consecutively (from partite sets different from $X_{\alpha}$) on its right up to $v_{k}$. Hence, there exists a vertex $v_{j}$ such that $v_{j}\notin X_{\alpha}$, $i<j<k$, and $v_{i}$ is not adjacent to $v_{j}$.
We consider the following cases:
  • Case 1. $v_{j}$ and $v_{k}$ belong to the same partite set $X_{\gamma}$. Then configuration (i) (Figure 8) occurs — a contradiction.
  • Case 2. $v_{i},v_{j},v_{k}$ belong to three distinct partite sets $X_{\alpha},X_{\beta},X_{\gamma}$ respectively.
    • Subcase 1: If $v_{j}$ is not adjacent to $v_{k}$, then configuration (ii) occurs — a contradiction.
    • Subcase 2: If $v_{j}$ is adjacent to $v_{k}$, then configuration (iii) occurs — a contradiction.
Thus, in every possible case, we reach a contradiction. Therefore, if $v_{i}$ and $v_{k}$ are adjacent, then $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{k}}\neq\emptyset$.
Hence, the constructed family of intervals $\mathcal{I}^{\prime}$ serves as an interval model of the $r$-partite graph $G$. Therefore, $G$ is an interval $r$-graph.
3 Conclusion
In this paper, we discussed various types of vertex orderings for the most generalized version of interval bigraphs, namely interval $r$-graphs. We also identified the corresponding forbidden patterns for this class with respect to certain specific orderings. For future research, one could aim to extend these concepts to the class of circular-arc graphs, with the goal of generalizing the notion of circular-arc bigraphs as well. It would also be interesting to investigate whether a finite set of forbidden patterns exists for these generalized classes.
Acknowledgement
The first author sincerely acknowledges the Council of Scientific and Industrial Research (CSIR), India, for providing financial support through the CSIR Fellowship (File No. 09/0028
(11986)/2021-EMR-I).

### v2

@@
 An $r$-partite graph is an interval $r$-graph if corresponding to each vertex we can assign an interval of the real line such that two vertices $u$ and $v$ of different partite sets are adjacent if and only if their corresponding intervals intersect.
-In this paper, we provide two vertex-ordering characterizations of interval $r$-graphs and identify forbidden patterns for interval $r$-graphs in terms of specific orderings of their vertices.
+In this paper, we provide two vertex-ordering characterizations of interval $r$-graphs, utilizing generalized interval orderings and $r$-interval orderings, and identify forbidden patterns for interval $r$-graphs in terms of specific orderings of their vertices.
 Keywords: interval $r$-graphs, vertex ordering, generalized interval ordering, $r$-interval ordering , forbidden patterns
@@
 The concept of interval bigraph was generalized by Brown [1].
-A graph $G=(V,E)$ is called interval $r$-graph if the vertex set $V$ can be partitioned into $r$ stable sets say $V_{1}$, $V_{2}$,…, $V_{r}$ and corresponding to each vertex we can assign an interval such that two vertices $u$ and $v$ are adjacent in $G$ if and only if the corresponding intervals $I_{u}$ and $I_{v}$ have non empty intersection and $u$ and $v$ belong to different partite set of vertices.
+A graph $G=(V,E)$ is called interval $r$-graph if the vertex set $V$ can be partitioned into $r$ stable sets say $V_{1}$, $V_{2}$,…, $V_{r}$ and corresponding to each vertex we can assign an interval such that two vertices $u$ and $v$ are adjacent in $G$ if and only if the corresponding intervals $I_{u}$ and $I_{v}$ have non-empty intersection and $u$ and $v$ belong to different partite sets of vertices.
 In this paper, we study different types of vertex orderings of interval $r$-graphs and, with the help of these orderings, we characterize the class of interval $r$-graphs.
@@
 We distinguish two cases according to the relative positions of $v_{i}$ and $v_{j}$ in the vertex ordering.
-Case 1: $i>j$.
+Case 1:
+When $i>j$.
 From the interval representation (see Figure 2), any vertex appearing between $v_{j}$ and $v_{i}$ in this ordering, and belonging to a partite set other than $X_{\beta}$, will be adjacent to $v_{j}$.
@@
 Figure 2: $I_{v_{i}}\cap I_{v_{j}}\neq\emptyset$, where $i>j$.
-Case 2: $i<j$.
-Similarly, any vertex appearing between $v_{i}$ and $v_{j}$ that does not belong to the same partite set $X_{\alpha}$ will be adjacent to $v_{i}$.
+Case 2:
+When $i<j$.
+Similarly, any vertex appearing between $v_{i}$ and $v_{j}$ that does not belong to the same partite set $X_{\alpha}$ will be adjacent to $v_{i}$ (see Figure 3).
 Hence, $R_{i}$ will contain the 1 corresponding to the entry $(i,j)$ in the adjacency matrix.
@@
 Therefore the class of intervals as constructed above is an interval model for the $r$-partite graph and so $G=(X_{1},X_{2},...,X_{r},E)$ is an interval $r$-graph.
-Example 1. Consider the following $3$-partite graph. Using the above theorem we show that it is an interval $3$-graph.
+Example 1. Consider the following $3$-partite graph in Figure 4. Using the above theorem we show that it is an interval $3$-graph.
 $v_{7}$$v_{5}$$v_{3}$$v_{4}$$v_{6}$$v_{10}$$v_{1}$$v_{2}$$v_{9}$$v_{8}$
@@
 […]
-Figure 5: The adjacency matrix of the 33 partite graph in Figure 2, where the rows and columns are arranged according to the increasing order of their indices, and corresponding $\mathcal{R}_{i}$’s and $\mathcal{C}_{j}$’s are shown in the elliptic regions.
-And its interval representation is as follows:
-Red: $I_{v_{1}}=[1,3]$, $I_{v_{6}}=[6,9]$, $I_{v_{10}}=[10,10]$
-Green: $I_{v_{2}}=[3,4]$, $I_{v_{8}}=[8,10]$, $I_{v_{9}}=[9,9]$
-Black: $I_{v_{3}}=[3,3]$, $I_{v_{4}}=[4,6]$, $I_{v_{5}}=[5,8]$, $I_{v_{7}}=[7,9]$.
+Figure 5: The adjacency matrix of the 3-partite graph in Figure 4, where the rows and columns are arranged according to the increasing order of their indices, and corresponding $\mathcal{R}_{i}$’s and $\mathcal{C}_{j}$’s are shown in the elliptic regions.
+Its interval representation is as follows: Red: $I_{v_{1}}=[1,3]$, $I_{v_{6}}=[6,9]$, $I_{v_{10}}=[10,10]$; Green: $I_{v_{2}}=[3,4]$, $I_{v_{8}}=[8,10]$, $I_{v_{9}}=[9,9]$; Black: $I_{v_{3}}=[3,3]$, $I_{v_{4}}=[4,6]$, $I_{v_{5}}=[5,8]$, $I_{v_{7}}=[7,9]$.
 Note that the zeros inside the elliptic regions of $\mathcal{R}_{5}$ and $\mathcal{C}_{5}$, as shown in Figure 5, do not belong to the sets $\mathcal{R}_{5}$ and $\mathcal{C}_{5}$.
@@
 […]
-Figure 8: Forbidden patterns.
+Figure 8: Forbidden patterns for interval $r$-graphs ($r\geq 3$) in terms of specific orderings of their vertices.
 Proof.
@@
 We consider the following cases:
-  • Case 1. $v_{j}$ and $v_{k}$ belong to the same partite set $X_{\gamma}$. Then configuration (i) (Figure 8) occurs — a contradiction.
+  • Case 1. $v_{j}$ and $v_{k}$ belong to the same partite set $X_{\gamma}$. Then configuration (i) (see Figure 8) occurs — a contradiction.
   • Case 2. $v_{i},v_{j},v_{k}$ belong to three distinct partite sets $X_{\alpha},X_{\beta},X_{\gamma}$ respectively.
-    • Subcase 1: If $v_{j}$ is not adjacent to $v_{k}$, then configuration (ii) occurs — a contradiction.
-    • Subcase 2: If $v_{j}$ is adjacent to $v_{k}$, then configuration (iii) occurs — a contradiction.
+    • Subcase 1: If $v_{j}$ is not adjacent to $v_{k}$, then configuration (ii) (see Figure 8) occurs — a contradiction.
+    • Subcase 2: If $v_{j}$ is adjacent to $v_{k}$, then configuration (iii) (see Figure 8) occurs — a contradiction.
 Thus, in every possible case, we reach a contradiction. Therefore, if $v_{i}$ and $v_{k}$ are adjacent, then $I^{\prime}_{v_{i}}\cap I^{\prime}_{v_{k}}\neq\emptyset$.
@@
 Acknowledgement
-The first author sincerely acknowledges the Council of Scientific and Industrial Research (CSIR), India, for providing financial support through the CSIR Fellowship (File No. 09/0028
-(11986)/2021-EMR-I).
+The first author sincerely acknowledges the Council of Scientific and Industrial Research (CSIR), India [File No.
+09/0028(11986)/2021-EMR-I], for providing financial support through a fellowship to pursue the Ph.D.
+degree.